OpenStudy (anonymous):

49/64 divided by 7/72

OpenStudy (anonymous):

answer is 63/8

Parth (parthkohli):

Yeah, I'm wrong

Parth (parthkohli):

I got the second last wrong

OpenStudy (anonymous):

7*7 / 8 * 8 / 7 / 8*9 = 7*7 * 8 * 9 / 8*8 * 7 = 7 *9 / 8 = 63 / 8

Parth (parthkohli):

= 49/64 * 72/7 = 7/64 * 72 = 7/8 * 9 = 63/8 = 7 7/8

Parth (parthkohli):

Guys, convert into mixed also.